Find the slope and the y intercept of 7x = 3y + 21
Find the slope if it exists of the line containing the pair of points (1,3) and (8,-2).
Find an equation of the line having the given slope and containing the given point m= 9, (7,6)
Find an equation of the line having the given slope and containing the given point m= 5/6, (4, -1).
Find an equation of the line containing the given pair of points. Express your answer in the form x=a, y=b, or y=mx + 6. (-2,-2) and (8,8).
Find an equation of the line containing the given pair of points. (-3, -9) and (-9, -2).

Answer by Alan3354(69443)

Find the slope and the y intercept of 7x = 3y + 21
Put it in slope-intercept form, which means solve for y.
3y = 7x - 21
y = (7/3)x - 7
Slope m = 7/3
y-intercept = -7, or the point (0,-7)
--------------------------------------
Find the slope if it exists of the line containing the pair of points (1,3) and (8,-2).
Slope, m, is diff in y over diff in x
m = (3-(-2))/(8-1)
m = 5/7
--------------
Find an equation of the line having the given slope and containing the given point m= 9, (7,6)
Use y-y1 = m*(x-x1) where (x1,y1) is the point.
y-6 = 9*(x-7)
y-6 = 9x-63
y = 9x - 57 (slope-intercept form)
9x - y = 57 (standard form)
-----------------------------
Find an equation of the line having the given slope and containing the given point m= 5/6, (4, -1)
Do it like the one above
-----------------------------
Find an equation of the line containing the given pair of points. Express your answer in the form x=a, y=b, or y=mx + 6. (-2,-2) and (8,8).
1st, find the slope as in the one above.
Then use either point, as I did above.
------------------------------
Find an equation of the line containing the given pair of points. (-3, -9) and (-9, -2)
Same again. If you have trouble, or want me to check your work, email me via the Thank you note.